# Multi-Mode Array Filtering of Resonance Fluorescence

### v3

@@ Abstract
We present a novel frequency-filtering method for measuring and calculating frequency-filtered photon-correlations. This novel method is a cavity-based system we call the multi-mode array filter, which consists of an array of tunable single-mode cavities that are equally spaced in frequency. By introducing a mode-dependent phase modulation, we produce a near rectangular frequency response, allowing us to increase the filter bandwidth – and thus the temporal response – without sacrificing frequency isolation. We model the frequency filtering using a cascaded quantum open systems approach which completely neglects any back-action of the filter onto the source system. This allows us to derive a closed set of operator moment equations for source and filter system operators, thus providing an extremely efficient method to calculate frequency-filtered first- and second-order correlation functions. We demonstrate this novel filtering method by applying it to a resonantly driven two-level atom. We present examples of frequency-filtered power spectra to demonstrate the improved frequency isolation of the multi-mode array filter over the single-mode filter. We then present results for the single-mode and multi-mode-array filtered second-order auto- and cross-correlation functions. These are compared against expressions derived in the secular approximation. The improved frequency isolation of the multi-mode array filter allows us to investigate new regimes of frequency-filtered photon correlations, such as two-photon leapfrog processes, and the effect of vanishing bandwidth on filtered auto-correlation functions.
@@ I Introduction
The majority of studies on atomic resonance fluorescence at this time were largely focused on the entire fluorescence spectrum of the atom. It was at the end of the 1970s when some of the first investigations were conducted on correlations between photons from different components of the Mollow triplet, as opposed to the whole spectrum [9, 10, 11, 12, 13], with the first experimental results reported by Aspect, Roger, Dalibard and Cohen-Tannoudji [14]. To achieve the separation of the two side-peaks, the atomic fluorescence was split into two channels, with each channel passing through an interference filter tuned to the respective target frequency, and modelled as a single-mode optical cavity [15, 16, 17]. Filtering such as this is essential to the measurement of frequency-resoled photon correlations as, typically, photon detectors are board bandwidth, and essentially non-selective in frequency.
@@
We begin in Section II by introducing the foundations of our filtering model, including a brief discussion of single-mode cavities and rectangular filters. We then present analytic and numerical results for the temporal and frequency response of the multi-mode array filter for different parameters. In Section III we review the dynamics and key phenomena of a resonantly driven two-level atom, which will serve as the source system with which to test the effectiveness of the filter. We then introduce the full atom-filter coupled system in Section IV and the moment equations relevant to solving for the two-time correlation functions. In Section V we present analytic results for frequency-filtered auto- and cross-correlations, comparing the multi-mode array-filtered results with those of a standard Lorentzian filter and approximate ideal solutions derived in the dressed-state basis. Finally, we conclude in Section VI with a summary of key results and a brief discussion on possible experimental implementations.
@@ II.3.2 Frequency response
$\bar{A}(\omega)\simeq\frac{-i\mathcal{E}_{d}e^{-i\omega t}}{\delta\omega\left(2N+1\right)}\mathrm{log}\left[\frac{\kappa-i\left(\omega-\omega_{N}\right)}{\kappa-i\left(\omega-\omega_{-N}\right)}\right],$ (23)
@@
×$\quad\quad\times\left|\mathrm{log}\left[\frac{\kappa+i\left(\omega-\omega_{N}\right)}{\kappa+i\left(\omega-\omega_{-N}\right)}\right]\right|^{2},$ (24)
@@ III.3.1 Auto-correlation functions
We can then use the quantum regression equations (see Refs. [45, 46] and [47], Sect. 1.5 for details) to derive second-order correlation functions for the central-, left-, and right-peak photon transitions, respectively:
@@ IV.1 Moment equation approach
With the master equation, Eq. (44), we can solve for the multi-mode array-filtered first- and second-order correlation functions using the quantum regression equations [45, 46]. For this filtering method, we allow for a large number of cavity modes, $N\gg 1$, each with an infinite dimension Hilbert space. Even by taking a conservative photon number truncation of up to two photons, the size of the total system’s density operator grows as $\left(2\times 3^{2N+1}\right)^{2}$. The computations then prove to be extremely demanding, and as such we were limited to $N=20$ on our current hardware.
@@
Figure 8: Frequency-filtered incoherent power spectrum of a single-mode (a, b), $N=0$, and multi-mode (c, d), $N=80$, filter. The filter resonance is resonant with the (a, c) central-peak, $\Delta\omega_{0}=0$, and (b, d) right-peak, $\Delta\omega_{0}=5\pi\gamma$, of the Mollow triplet, with halfwidths $K=2\gamma$ (blue, solid), $K=4\gamma$ (orange, dashed), $K=16\gamma$ (green, dot-dash). The filtered spectra are compared against the unfiltered Mollow triplet (grey, dotted). Other parameters are $\Omega=5\pi\gamma,m=1$ and, for the multi-mode array filter $\delta\omega=K/N,\kappa=2.5\delta\omega$.
@@ IV.2 First-order correlation function
While we do not give explicit expressions for the steady state expectation values in Eqs. (55-56), we can compute these numerically from the moment equations given in Appendix A.
Using Eq. (IV.2), we compute the incoherent power spectrum of the single- and multi-mode array filters for increasing halfwidths, shown in Fig. 8. For the multi-mode array filter, we choose a large number of modes, $N=80$, to ensure a sharp cut-off in the frequency response. Comparing the single-mode filter, Figs. 8(a, b), to the unfiltered Mollow triplet (grey dotted line), we see two distinct features. Firstly, for the smallest halfwidth shown, $K=2\gamma$, the output resonance is much narrower than the target peak. Increasing the halfwidth worsens the frequency isolation, as the Lorentzian response overlaps with the other peaks of the Mollow triplet.
Figures 8(c, d) show a dramatic drop-off in the incoherent power spectra of the multi-mode array filtered fluorescence. As with the single-mode filter, for the smallest halfwidth, $K=2\gamma$, the linewidth of the filtered fluorescence is much narrower than the natural linewidth. With the largest halfwidth, $K=16\gamma$, the multi-mode array filter performs similarly to the single-mode filter, as the halfwidth encompasses both side-peaks of the Mollow triplet. The intermediate halfwidth, at $K=8\gamma$, however, demonstrates the sharp frequency response cut-off of the multi-mode array filter. This gives a good indication of the optimal halfwidth regime for accurately measured frequency-filtered photon correlations.
@@ V Frequency-Filtered Photon Correlations
We note here the difference between Eq. (59c) and Eq. (43) is the factor of $1/2$ to account for the 50:50 beam splitter guiding the fluorescence into the two multi-mode array filters. For this two-filter system, we wish to correlate photons emitted from filter B some time $\tau$ after a detection of an emission from mode A. We therefore introduce the frequency-filtered second-order cross-correlation function:
Figure 9: Frequency-filtered auto-correlation function of a single-mode (a, b), $N=0$, and multi-mode (c, d), $N=80$, filter. The filter resonance is resonant with the (a, c) central-peak, $\Delta\omega_{0}=0$, and (b, d) right-peak, $\Delta\omega_{0}=5\pi\gamma$, of the Mollow triplet, with halfwidths $K=2\gamma$ (blue, solid), $K=4\gamma$ (orange, dashed), $K=8\gamma$ (green, dot-dash). The filtered correlation functions are plotted against the secular approximation equations, (a, b) Eq. (38a) and (c, d) Eq. (38c). Other parameters are $\Omega=5\pi\gamma,m=1$ and, for the multi-mode array filter $\delta\omega=K/N,\kappa=2.5\delta\omega$.
@@
We again use the moment equation approach as mentioned in the previous section, with the moment equations given in Appendix A.
@@ V.1 Auto-correlations
There are four parameters of the multi-mode array filter: the number of modes, $N$; the frequency spacing between each mode, $\delta\omega$; the halfwidth of each individual mode, $\kappa$; and, the size of the phase modulation, $m$. For this work, however, we will fix the number of modes to $N=80$, and set the frequency spacing based on the effective halfwidth, $K=N\delta\omega$. As established in Section II, we will also assume the halfwidth of each mode to be much smaller than the effective halfwidth, with $\kappa=\frac{\delta\omega}{4}=\frac{K}{4N}$. We will also set the size of the phase modulation to $m=1$. For this paper we will only consider the effect that the effective halfwidth, $K$, has on the single- and multi-mode-filtered correlation functions. For a detailed analysis on the effect of other parameters, see Ref. [48], Sect. 6.3.1.
@@ V.1.1 Varying the halfwidth
In Fig. 9 we depict sets of frequency-filtered photon correlation functions for increasing halfwidths – $K=2\gamma$ (blue solid line), $4\gamma$ (orange dashed line), and $8\gamma$ (green dot-dash line) – for both the single-mode filter, Figs. 9(a, b), and the multi-mode array filter, Figs. 9(c, d). We compare the frequency-filtered second-order correlation functions to the ideal correlation functions derived in the secular approximation (grey dotted line), Eqs. (38a) and (38c).
@@
In the bad cavity limit, as $K/\gamma\rightarrow\infty$, the “filtered” field is indistinguishable from the unfiltered atomic fluorescence; the photon correlations exhibit large Rabi oscillations, of frequency $\Omega$, as the atomic state oscillates between the ground and excited states. For the single-mode filter, Figs. 9(a, b), these Rabi oscillations are visible even at the smaller halfwidth, $K=2\gamma$. As the halfwidth is increased, these oscillations get larger; indicating, in the temporal domain, that the frequency isolation is worsening.
@@
There is an intermediate region – $K\sim 2.5\gamma$ to $\sim 8\gamma$ – where the initial correlation value approaches the value expected from the secular approximation: $g^{(2)}(0,0;0,0)=1$ for the central peak, Fig. 10(a); and $g^{(2)}(\Omega,0;\Omega,0)=0$ for the right peak, Fig. 10(b). At first glance, in Fig. 10, it seems that the single-mode filter would perform better than the multi-mode array filter, as there is a direct match with the initial auto-correlation value in the secular approximation at $K\sim 5.5\gamma$. We have seen, however, in Figs. 8 and 9 that this halfwidth is still too large to effectively isolate the central peak.
@@ V.1.2 Filtering within the peak
Focusing first on the single-mode filtered-correlations, we see similar regimes for both the filtered central and right-peaks (Fig. 12a and b, respectively). As previously mentioned, very large filter halfwidths, $K\gg\Omega$, have a negligible effect on the atomic fluorescence, resulting in near perfect antibunching. As the halfwidth decreases to the range $\gamma<K<\Omega$, the filter mostly rejects the non-target peaks, resulting in auto-correlations close to the secular approximation, as discussed in the previous section.
Figure 12: Initial value of the correlation function of the (a) central, $\Delta\omega_{0}=0$ and (b) right peak, $\Delta\omega_{0}=5\pi\gamma$, for the multi-mode (blue, solid), $N=80$, and single-mode (orange, dashed), $N=0$, filters. Other parameters are $\Omega=5\pi\gamma,\delta\omega=K/N$, and $\kappa=2.5\delta\omega$.
@@
The multi-mode array filter produces similar results to the single-mode filter: large halfwidths give near perfect antibunching, intermediate halfwidths can isolate the target peaks, and vanishingly small halfwidths result in thermalisation for the filtered right peak and second-order coherence for the filtered central peak. We do see some slight differences, such as a shift in the bunching peak for the filtered central peak, due to the improved frequency isolation. The most striking difference is the emergence of peaks at $K\sim\Omega$ for both the filtered central- and right-peaks, and a second peak at $K\sim 2\Omega$ for the filtered right-peak. Similar results can be seen in a perfect rectangular filter (see Figs. 4 and 5 of Ref. [36]). Kamide et al. suggest that these peaks are due to constructive interference in the cascaded decay channels.
@@ V.2.1 Temporal evolution of peak-to-peak correlations
We note here that we have chosen different halfwidths for the single- and multi-mode array filters, with the same numerical optimisation method as in Fig. 11, hence why the expected short-time behaviours differ. With a halfwidth roughly three times larger, however, the visible Rabi oscillations in the multi-mode-filtered cross-correlations are still significantly reduced when compared to the single-mode filter. Larger $K$ for the single-mode filter would result in much larger Rabi oscillations in the cross-correlations.
@@ Appendix A Moment Equations for the Two-Filter System
Routines written in Fortan90 compute the steady states of these moment equations using the LAPACK matrix inversion routines [59]. The routines can be found at the following GitHub repository: github.com/jnga773/multi-mode-filter.

### v4

@@ Abstract
We present a frequency-filtering method for measuring and calculating frequency-filtered photon-correlations. This method is a cavity-based system we call the multi-mode array filter, which consists of an array of tunable single-mode cavities that are equally spaced in frequency. By introducing a mode-dependent phase modulation, we produce a near rectangular frequency response, allowing us to increase the filter bandwidth – and thus the temporal response – without sacrificing frequency isolation. We model the frequency filtering using a cascaded quantum open systems approach which completely neglects any back-action of the filter onto the source system. This allows us to derive a closed set of operator moment equations for source and filter system operators, thus providing an extremely efficient method to calculate frequency-filtered first- and second-order correlation functions. We demonstrate this filtering method by applying it to a resonantly driven two-level atom. We present examples of frequency-filtered power spectra to demonstrate the improved frequency isolation of the multi-mode array filter over the single-mode filter. We then present results for the single-mode and multi-mode-array filtered second-order auto- and cross-correlation functions. These are compared against expressions derived in the secular approximation. The improved frequency isolation of the multi-mode array filter allows us to investigate new regimes of frequency-filtered photon correlations, such as two-photon leapfrog processes, and the effect of vanishing bandwidth on filtered auto-correlation functions.
@@ I Introduction
The majority of studies on atomic resonance fluorescence at this time were largely focused on the entire fluorescence spectrum of the atom. It was at the end of the 1970s when some of the first investigations were conducted on correlations between photons from different components of the Mollow triplet, as opposed to the whole spectrum [9, 10, 11, 12, 13], with the first experimental results reported by Aspect, Roger, Dalibard and Cohen-Tannoudji [14]. To achieve the separation of the two side-peaks, the atomic fluorescence was split into two channels, with each channel passing through an interference filter tuned to the respective target frequency, and modelled as a single-mode optical cavity [15, 16, 17]. Filtering such as this is essential to the measurement of frequency-resoled photon correlations as, typically, photon detectors are broad bandwidth, and essentially non-selective in frequency.
@@
We begin in Sect. II by introducing the foundations of our filtering model, including a brief discussion of single-mode cavities and rectangular filters. We then present analytic and numerical results for the temporal and frequency response of the multi-mode array filter for different parameters. In Sect. III we review the dynamics and key phenomena of a resonantly driven two-level atom, which will serve as the source system with which to test the effectiveness of the filter. We then introduce the full atom-filter coupled system in Sect. IV and the moment equations relevant to solving for the two-time correlation functions. In Sect. V we present analytic results for frequency-filtered auto- and cross-correlations, comparing the multi-mode array-filtered results with those of a standard Lorentzian filter and approximate ideal solutions derived in the dressed-state basis. Finally, we conclude in Sect. VI with a summary of key results and a brief discussion on possible experimental implementations.
@@ II.3.2 Frequency response
$\bar{A}(\omega)\simeq\frac{-i\mathcal{E}_{d}e^{-i\omega t}}{\delta\omega\left(2N+1\right)}\mathrm{ln}\left[\frac{\kappa-i\left(\omega-\omega_{N}\right)}{\kappa-i\left(\omega-\omega_{-N}\right)}\right],$ (23)
@@
×$\quad\quad\times\left|\mathrm{ln}\left[\frac{\kappa+i\left(\omega-\omega_{N}\right)}{\kappa+i\left(\omega-\omega_{-N}\right)}\right]\right|^{2},$ (24)
@@ III.3.1 Auto-correlation functions
We can then use the quantum regression equations (see Refs. [45, 46] and Sec. 1.5 of [47] for details) to derive second-order correlation functions for the central-, left-, and right-peak photon transitions, respectively:
@@ IV.1 Moment equation approach
With the master equation, Eq. (44), we can solve for the multi-mode array-filtered first- and second-order correlation functions using the quantum regression equations [45, 46]. For this filtering method, we allow for a large number of cavity modes, $N\gg 1$, each with an infinite dimension Hilbert space. Even by taking a conservative photon number truncation of up to two photons, the size of the total system’s density operator grows as $\left(2\times 3^{2N+1}\right)^{2}$. The computations then prove to be extremely demanding, and as a result we were limited to $N=20$ on our current hardware.
@@
Figure 8: Frequency-filtered incoherent power spectrum of a single-mode (a, b), $N=0$, and multi-mode (c, d), $N=80$, filter. The filter resonance is resonant with the (a, c) central-peak, $\Delta\omega_{0}=0$, and the (b, d) right-peak, $\Delta\omega_{0}=5\pi\gamma$, of the Mollow triplet, with halfwidths $K=2\gamma$ (blue, solid), and $K=4\gamma$ (orange, dashed), $K=16\gamma$ (green, dot-dash). The filtered spectra are compared against the unfiltered Mollow triplet (grey, dotted). Other parameters are $\Omega=5\pi\gamma,m=1$ and, for the multi-mode array filter $\delta\omega=K/N,\kappa=2.5\delta\omega$.
@@ IV.2 First-order correlation function
While we do not give explicit expressions for the steady state expectation values in Eqs. (55) and (56), we can compute these numerically from the moment equations given in the Appendix.
Using Eq. (IV.2), we compute the incoherent power spectrum of the single- and multi-mode array filters for increasing halfwidths, shown in Fig. 8. For the multi-mode array filter, we choose a large number of modes, $N=80$, to ensure a sharp cut-off in the frequency response. Comparing the single-mode filter, Figs. 8(a) and 8(b), to the unfiltered Mollow triplet (grey dotted line), we see two distinct features. Firstly, for the smallest halfwidth shown, $K=2\gamma$, the output resonance is much narrower than the target peak. Increasing the halfwidth worsens the frequency isolation, as the Lorentzian response overlaps with the other peaks of the Mollow triplet.
Figures 8(c) and 8(d) show a dramatic drop-off in the incoherent power spectra of the multi-mode array filtered fluorescence. As with the single-mode filter, for the smallest halfwidth, $K=2\gamma$, the linewidth of the filtered fluorescence is much narrower than the natural linewidth. With the largest halfwidth, $K=16\gamma$, the multi-mode array filter performs similarly to the single-mode filter, as the halfwidth encompasses both side-peaks of the Mollow triplet. The intermediate halfwidth, at $K=8\gamma$, however, demonstrates the sharp frequency response cut-off of the multi-mode array filter. This gives a good indication of the optimal halfwidth regime for accurately measured frequency-filtered photon correlations.
@@ V Frequency-Filtered Photon Correlations
We note here the difference between Eqs. (59c) and (43) is the factor of $1/2$ to account for the 50:50 beam splitter guiding the fluorescence into the two multi-mode array filters. For this two-filter system, we wish to correlate photons emitted from filter B some time $\tau$ after a detection of an emission from mode A. We therefore introduce the frequency-filtered second-order cross-correlation function:
Figure 9: Frequency-filtered auto-correlation function of a single-mode (a, b), $N=0$, and multi-mode (c, d), $N=80$, filter. The filter resonance is resonant with the (a, c) central-peak, $\Delta\omega_{0}=0$, and (b, d) right-peak, $\Delta\omega_{0}=5\pi\gamma$, of the Mollow triplet, with halfwidths $K=2\gamma$ (blue, solid), $K=4\gamma$ (orange, dashed), and $K=8\gamma$ (green, dot-dash). The filtered correlation functions are plotted against the secular approximation equations, (a, b) Eq. (38a) and (c, d) Eq. (38c). Other parameters are $\Omega=5\pi\gamma,m=1$ and, for the multi-mode array filter $\delta\omega=K/N,\kappa=2.5\delta\omega$.
@@
We again use the moment equation approach as mentioned in the previous section, with the moment equations given the Appendix.
@@ V.1 Auto-correlations
There are four parameters of the multi-mode array filter: the number of modes, $N$; the frequency spacing between each mode, $\delta\omega$; the halfwidth of each individual mode, $\kappa$; and, the size of the phase modulation, $m$. For this work, however, we will fix the number of modes to $N=80$, and set the frequency spacing based on the effective halfwidth, $K=N\delta\omega$. As established in Section II, we will also assume the halfwidth of each mode to be much smaller than the effective halfwidth, with $\kappa=\frac{\delta\omega}{4}=\frac{K}{4N}$. We will also set the size of the phase modulation to $m=1$. For this paper we will only consider the effect that the effective halfwidth, $K$, has on the single- and multi-mode-filtered correlation functions. For a detailed analysis on the effect of other parameters, see Ref. [48], Sec. 6.3.1.
@@ V.1.1 Varying the halfwidth
In Fig. 9 we depict sets of frequency-filtered photon correlation functions for increasing halfwidths – $K=2\gamma$ (blue solid line), $4\gamma$ (orange dashed line), and $8\gamma$ (green dot-dash line) – for both the single-mode filter, Figs. 9(a) and 9(b), and the multi-mode array filter, Figs. 9(c) 9(d). We compare the frequency-filtered second-order correlation functions to the ideal correlation functions derived in the secular approximation (grey dotted line), Eqs. (38a) and (38c).
@@
In the bad cavity limit, as $K/\gamma\rightarrow\infty$, the “filtered” field is indistinguishable from the unfiltered atomic fluorescence; the photon correlations exhibit large Rabi oscillations, of frequency $\Omega$, as the atomic state oscillates between the ground and excited states. For the single-mode filter, Figs. 9(a) and 9(b), these Rabi oscillations are visible even at the smaller halfwidth, $K=2\gamma$. As the halfwidth is increased, these oscillations get larger; indicating, in the temporal domain, that the frequency isolation is worsening.
@@
There is an intermediate region, $K\sim 2.5\gamma$ to $\sim 8\gamma$, where the initial correlation value approaches the value expected from the secular approximation: $g^{(2)}(0,0;0,0)=1$ for the central peak, Fig. 10(a); and $g^{(2)}(\Omega,0;\Omega,0)=0$ for the right peak, Fig. 10(b). At first glance, in Fig. 10, it seems that the single-mode filter would perform better than the multi-mode array filter, as there is a direct match with the initial auto-correlation value in the secular approximation at $K\sim 5.5\gamma$. We have seen, however, in Figs. 8 and 9 that this halfwidth is still too large to effectively isolate the central peak.
@@ V.1.2 Filtering within the peak
Focusing first on the single-mode filtered-correlations, we see similar regimes for both the filtered central and right-peaks [Figs. 12(a) and 12(b), respectively]. As previously mentioned, very large filter halfwidths, $K\gg\Omega$, have a negligible effect on the atomic fluorescence, resulting in near perfect antibunching. As the halfwidth decreases to the range $\gamma<K<\Omega$, the filter mostly rejects the non-target peaks, resulting in auto-correlations close to the secular approximation, as discussed in the previous section.
Figure 12: Initial value of the correlation function of the (a) central peak, $\Delta\omega_{0}=0$ and (b) right peak, $\Delta\omega_{0}=5\pi\gamma$, for the multi-mode (blue, solid), $N=80$, and single-mode (orange, dashed), $N=0$, filters. Other parameters are $\Omega=5\pi\gamma,\delta\omega=K/N$, and $\kappa=2.5\delta\omega$.
@@
The multi-mode array filter produces results similar to the single-mode filter: large halfwidths give near perfect antibunching, intermediate halfwidths can isolate the target peaks, and vanishingly small halfwidths result in thermalisation for the filtered right peak and second-order coherence for the filtered central peak. We do see some slight differences, such as a shift in the bunching peak for the filtered central peak, due to the improved frequency isolation. The most striking difference is the emergence of peaks at $K\sim\Omega$ for both the filtered central- and right-peaks, and a second peak at $K\sim 2\Omega$ for the filtered right-peak. Similar results can be seen in a perfect rectangular filter (see Figs. 4 and 5 of Ref. [36]). Kamide et al. suggest that these peaks are due to constructive interference in the cascaded decay channels.
@@ V.2.1 Temporal evolution of peak-to-peak correlations
We note here that we have chosen different halfwidths for the single- and multi-mode array filters, with the same numerical optimisation method as in Fig. 11, which is why the expected short-time behaviours differ. With a halfwidth roughly three times larger, however, the visible Rabi oscillations in the multi-mode-filtered cross-correlations are still significantly reduced when compared to the single-mode filter. Larger $K$ for the single-mode filter would result in much larger Rabi oscillations in the cross-correlations.
@@ Appendix A Moment Equations for the Two-Filter System
Routines written in Fortan90 compute the steady states of these moment equations using the LAPACK matrix inversion routines [59]. The routines can be found in a GitHub repository [60].